OpenStudy (anonymous):

using the foil method: FOIL. (a+4)(a-2)

OpenStudy (anonymous):

i know it means, first, outer, inner, last. BUT i multiply them? so a times a? first = a^2?

OpenStudy (amistre64):

consider the following: 2*2 = 2^2 3*3 = 3^2 5*5 = 5^2 a*a = ?

OpenStudy (anonymous):

a^2

OpenStudy (amistre64):

very good :)

OpenStudy (anonymous):

I hate foil. Such a waste of brain space ;p

OpenStudy (anonymous):

Yes, you multiply them.

OpenStudy (anonymous):

so : a^2+2a-8 is the answer? did i do it right? (simplified, of course)

OpenStudy (anonymous):

That is correct!

OpenStudy (anonymous):

foil is not a favorite, either is most of the other math...haha

OpenStudy (anonymous):

Well my issue with foil is that it only works with products in the form: (a+b)(c+d) But you already learned distribution which is much more robust, and is the underlying reason why foil works, so why not just distribute.

OpenStudy (anonymous):

x(4x-5)(x+8) = 4x^2-4x+8 ?

OpenStudy (amistre64):

distributing is the proper ... way to go about it.

OpenStudy (anonymous):

not quite lish.

OpenStudy (amistre64):

id multiply the last two "terms" first then multiply in that x

OpenStudy (anonymous):

Start by just leaving the first x factor out in front and just foil the back two, then distribute the x into what you got from foil

OpenStudy (anonymous):

@Amistre s/term/factor/

OpenStudy (anonymous):

<3

OpenStudy (amistre64):

:) its me florida accent lol

OpenStudy (anonymous):

I'm a pedant. Sorry again. =)

OpenStudy (anonymous):

4x^3-27x^2-40x ?

OpenStudy (anonymous):

yes. Nice job.

OpenStudy (anonymous):

wooo hoooo. i thank you guys for putting up with all my questions!

OpenStudy (anonymous):

It's what we do =)

OpenStudy (anonymous):

when i use the foil for this (2s+6)(2s-6) when i get my answer 4s^2+s-36 should that s not be in there since the product of those inner numbers is zero?